# Robust Polyp Detection and Diagnosis through Compositional Prompt-Guided Diffusion Models

### v1

@@ Abstract
Keywords: Diffusion models, compositional prompts, medical image generation, polyp detection and diagnosis.
@@ 1 Introduction
Efforts to improve generalization focus on collecting diverse, well-annotated datasets, which are critical for reliable deep learning performance. For instance, PolypGen [7] captures polyp diversity by combining data from six centers. However, creating such datasets is costly, labor-intensive, and constrained by privacy concerns, limiting scalability [8]. Traditional data augmentation techniques, such as affine transformations (e.g., rotation, flipping, cropping) and color modifications, have been widely used to expand datasets and improve generalization. However, these methods offer limited variability and fail to capture the complexity of medical images or generate novel visual features [9]. To address these limitations, synthetic data generation using models like Generative Adversarial Networks (GANs) [10] has emerged as a solution, increasing dataset variability and enhancing generalization. In medical imaging, this approach generates diverse and realistic representations, improving model performance across clinical scenarios. However, GANs often struggle with consistent quality and diversity, limiting their practical effectiveness [11]. Diffusion models (DMs) provide a promising alternative, refining noise into structured data to generate stable and diverse images. DMs address domain generalization challenges and enable targeted applications through conditional prompts [12], improving performance in tasks like skin lesion classification [13] and colonoscopy analysis [14].
@@
  • Colorectal Expert-Annotated Dataset: In collaboration with colorectal endoscopists, we provide text annotations for open-source datasets, including reports on polyp size, type, surface characteristics, and pathology. This dataset serves as a valuable supplement to existing public datasets, enriching the breadth of colorectal research.
@@ 3 Methods
Figure 2: The framework integrates diverse medical annotations into compositional prompts, processed via the Prompt Spectrum to capture both spatial and clinical details, as shown in (a) (Section 3.2). During denoising, a conditional U-Net progressively refines latent variables from low- to high-frequency components, as shown in (b) (Section 3.3).
@@ 3.1 Model Overview
Our framework generates high-quality synthetic polyp images using compositional prompts derived from diverse annotations, including segmentation masks, bounding boxes, textual descriptions, and vague class labels, as illustrated in Fig. 2. To integrate these prompts, we propose the Prompt Spectrum (Section 3.2), which categorizes the embeddings $\mathbf{P}$ into coarse and fine components to capture both structural layouts and detailed clinical features. These embeddings condition the diffusion model $\mathcal{D}_{\theta}$ during the Prompt-guided Diffusion Process (Section 3.3), enabling a progressive integration of spatial and semantic details to balance global structures and intricate clinical information.
@@ 3.2.3 Prompt for Mask and Bounding Box Annotations
The process begins by converting the binary mask to a one-hot representation, distinctly defining foreground and background regions. The one-hot mask is then downsampled through interpolation stages, reducing resolution while preserving key spatial characteristics. A final linear layer produces a compact mask embedding, $\mathbf{p}_{\text{mask}}$, which encodes both coarse spatial layout and fine boundary details, leveraging the mask’s inherent structural information.
For bounding boxes, a similar approach is applied. We generate random elliptical masks based on bounding box dimensions, capturing overall spatial structure. These masks are processed by the mask_encoder to produce layout embeddings, $\mathbf{p}_{\text{layout}}$, that represent the bounding box’s size and position without fine-grained detail.
@@ 3.2.4 Prompt Learning for Vague Class Annotations
In scenarios where class labels are ambiguous or vague, such as when the annotation is not a specific mask or text but rather a reference image group representing a vague class, we adopt a Prompt Learning approach integrated within the Prompt Spectrum. This approach utilizes specialized probabilistic embeddings that allow the model to handle uncertainty by generating prompts that represent a distribution of possible outcomes, rather than a single deterministic label.
@@ 3.3 Prompt-guided Diffusion Process
In this work, building on compositional prompts and the Prompt Spectrum, we extend the Stable Diffusion framework [28] into the Progressive Spectrum Diffusion Model (PSDM). PSDM guides the image generation process through multi-granularity prompts $\mathbf{p}_{i}$, ensuring alignment with complex medical annotation requirements.
@@ 3.4 Training Strategy
Generating accurate images from diverse prompts while learning sequentially from different annotation types introduces a significant challenge known as catastrophic forgetting. This occurs when, as new annotations are introduced, the model ”forgets” how to generate images based on earlier annotations, diminishing its ability to retain previously learned information [30, 17].
@@
[…]
@@ 4.1.2 Downstream Tasks
Polyp Segmentation: For the segmentation task, following the PraNet standard, we followed the PraNet standard and trained our model on 1,450 image-mask pairs sourced from the Kvasir and CVC-ClinicDB datasets. Additionally, the PolypGen dataset (centers C1-C6) was incorporated into the evaluation process to ensure a comprehensive assessment across diverse clinical environments. To improve the model’s generalization, we introduced two augmented datasets: the Segmentation Mask Only set, containing 1,450 image-mask pairs generated solely from the original training masks, and the Segmentation Mask-Text Description set, with 1,450 pairs generated using both textual descriptions and segmentation masks from the same training set.
@@
Polyp Detection: For detection, we used the PolypGen dataset, excluding 64 images from C3 without positive labels. The training set included images from centers C1 to C5 and sequences Seq1-15 (Positive) and Seq1-13 (Negative). The test set comprised images from C6 and sequences Seq16-23 (Positive) and Seq14-23 (Negative). We also introduced two augmented datasets for detection: the Bounding Box Only set, with bounding boxes regressed from the Segmentation Mask Only set, and the Bounding Box-Text Description set, with bounding boxes from the Segmentation Mask-Text Description set. Bounding boxes were defined by the smallest rectangle enclosing each polyp in the segmentation masks.
@@ 4.1.3 Baseline Models and Implementations
For the segmentation task, we employ three state-of-the-art models: PraNet [41], Polyp-PVT [42], and Polyp-CASCADE [43]. PraNet was selected for its proven effectiveness in capturing both local and global features, Polyp-PVT for its superior performance in handling multi-scale features, and Polyp-CASCADE for its robust architecture in medical segmentation tasks. For the classification task, we adopt ResNet [44], a widely used model in medical image analysis known for its effectiveness in classification tasks. For the polyp detection task, we utilize YOLOv5 [45], chosen for its high performance and efficiency in real-time object detection, which is critical in clinical settings. The hyperparameters for each model are provided in Table 2, categorized by task.
@@
Table 3: mDice and mIoU results for standard segmentation datasets. Best results are bold, and second-best are underlined
@@
$\text{PSDM}^{\text{S}}$ 88.37 81.31 92.30 87.60 89.76 84.12 74.00 66.45 63.55 56.95 75.91 69.09
$\text{PSDM}^{\text{M}}$ 88.49 81.64 92.08 86.99 89.37 83.72 73.57 66.21 67.81 60.98 76.70 69.89
@@
$\text{PSDM}^{\text{S}}$ 90.74 84.33 93.03 88.14 92.09 87.40 80.38 72.50 81.22 73.13 83.82 76.63
$\text{PSDM}^{\text{M}}$ 89.13 82.16 93.28 88.39 91.21 86.12 81.23 73.05 78.73 70.58 83.40 75.96
@@
$\text{PSDM}^{\text{S}}$ 88.74 82.01 93.50 88.92 92.34 87.66 80.86 72.88 79.33 71.45 83.50 76.31
$\text{PSDM}^{\text{M}}$ 89.26 82.64 94.02 89.44 92.01 87.16 82.28 73.93 80.31 72.51 84.45 77.10
Table 4: mDice and mIoU results for PolypGen. The best results are bold, and the second-best results are underlined
@@
$\text{PSDM}^{\text{S}}$ 80.75 73.92 74.05 68.31 87.51 81.06 37.80 32.52 58.03 50.08 76.07 70.01 71.76 65.38
$\text{PSDM}^{\text{M}}$ 81.89 75.01 73.60 67.92 87.35 80.78 39.52 34.26 60.80 52.54 76.70 70.25 72.48 66.01
@@
$\text{PSDM}^{\text{S}}$ 85.11 77.57 75.84 69.97 88.76 82.00 44.28 38.16 67.11 57.55 76.56 70.64 75.42 68.47
$\text{PSDM}^{\text{M}}$ 84.77 77.60 76.23 70.69 88.93 82.41 43.86 37.73 68.03 58.71 76.69 70.80 75.56 68.84
@@
$\text{PSDM}^{\text{S}}$ 85.77 79.12 76.17 70.85 90.36 84.37 44.35 38.67 67.13 58.59 78.23 72.65 76.18 69.94
$\text{PSDM}^{\text{M}}$ 86.17 79.68 76.37 71.11 89.71 83.64 45.25 39.94 68.04 58.88 77.67 72.02 76.32 70.06
@@ 4.2.1 Polyp Segmentation
In our experiments, we evaluated the performance of the segmentation models using two key metrics: mean Dice coefficient (mDice) and mean Intersection over Union (mIoU). These metrics are crucial in medical image segmentation, as they measure the overlap between the predicted segmentations and the ground truth, with higher values indicating better segmentation performance.
@@
Overall, these findings suggest that the proposed augmentation strategies consistently enhance models’ ability to generalize across both standard and challenging datasets. By integrating spatial and textual information through compositional prompts, the models show a superior capacity to handle variability in polyp characteristics. This improved generalization is crucial for real-world medical applications, where polyps exhibit significant heterogeneity in appearance, and accurate segmentation plays vital role in diagnostic precision.

### v2

# IEEEexample:BSTcontrol

@@ Abstract
{IEEEkeywords}
Diffusion models, compositional prompts, medical image generation, polyp detection and diagnosis.
@@ 1 Introduction
Efforts to improve generalization focus on collecting diverse, well-annotated datasets, which are critical for reliable deep learning performance. For instance, PolypGen [7] captures polyp diversity by combining data from six centers. However, creating such datasets is costly, labor-intensive, and constrained by privacy concerns, limiting scalability [8]. Traditional data augmentation techniques, such as affine transformations (e.g., rotation, flipping, cropping) and color modifications, have been widely used to expand datasets and improve generalization. However, these methods offer limited variability and fail to capture the complexity of medical images or generate novel visual features [9]. To address these limitations, synthetic data generation using models like Generative Adversarial Networks (GANs) [10] has emerged as a solution, increasing dataset variability and enhancing generalization. In medical imaging, this approach generates diverse and realistic representations, improving model performance across clinical scenarios. However, GANs often struggle with consistent quality and diversity, limiting their practical effectiveness [11]. Diffusion models (DMs) provide a promising alternative, refining noise into structured data to generate stable and diverse images. DMs address domain generalization challenges and enable targeted applications through conditional prompts [12], improving performance in tasks like skin lesion classification [13] and colonoscopy analysis [14][15].
@@
  • Colorectal Expert-Annotated Dataset: In collaboration with colorectal endoscopists, we provide text annotations for open-source datasets, including reports on polyp size, type, surface characteristics, and pathology. All annotations were created by a expert endoscopist, with ambiguous cases further reviewed and reconciled by a second senior clinician to ensure consistency. This dataset enriches existing public resources for colorectal research.
@@ 3 Methods
Figure 2: The framework integrates diverse annotations into compositional prompts via the Prompt Spectrum to capture spatial and clinical details (a, Sec. 3.2). During denoising, a conditional U-Net refines latent variables from low to high frequencies, with arrow thickness indicating each component’s contribution (b, Sec. 3.3).
@@ 3.1 Model Overview
Our framework generates high-quality synthetic polyp images by strategically combining compositional prompts derived from diverse annotations, including segmentation masks, bounding boxes, textual descriptions, and prototype images, as illustrated in Fig. 2. Each annotation type is categorized into coarse or fine prompts based on spatial detail and semantic precision, enabling joint control over both localization and morphological realism. Specifically, segmentation masks and bounding boxes constitute coarse prompts, primarily encoding spatial layout and broad localization without intricate appearance details. Textual reports include both coarse spatial characteristics (e.g., polyp size and location) and fine-grained semantic features (e.g., texture, color, polyp type, and pathology). Prototype images specifically address fine-level attributes, especially for rare or uncertain polyp classes with limited detailed annotations. To integrate these prompts, we propose the Prompt Spectrum (Section 3.2), which categorizes the embeddings $\mathbf{P}$ into coarse and fine components to capture both structural layouts and detailed clinical features. These embeddings condition the diffusion model $\mathcal{D}_{\theta}$ during the Prompt-guided Diffusion Process (Section 3.3), enabling a progressive integration of spatial and semantic details to balance global structures and intricate clinical information.
@@ 3.2.3 Prompt for Mask and Bounding Box Annotations
The process begins by converting the binary mask to a one-hot representation, distinctly defining foreground and background regions. The one-hot mask is then downsampled through interpolation stages, reducing resolution while preserving key spatial characteristics. A final linear layer produces a compact mask embedding, $\mathbf{p}_{\text{mask}}$.
For bounding boxes, a similar approach is applied. We generate random elliptical masks based on bounding box dimensions, capturing overall spatial structure. Specifically, to integrate bounding boxes into our prompt spectrum and support compositional generation without requiring additional encoder training, we approximate coarse spatial layouts by generating an elliptical mask that precisely fits (i.e., tangent to) each bounding box. We then apply slight perturbations along the ellipse boundary to better simulate realistic polyp morphologies observed clinically. These masks are processed by the mask_encoder to produce layout embeddings, $\mathbf{p}_{\text{layout}}$, that represent the bounding box’s size and position without fine-grained detail.
@@ 3.2.4 Prompt Learning for Vague Class Annotations
In scenarios where class labels are ambiguous or vague, such as when the annotation is not a specific mask or text but rather a prototype image group representing a vague class, we adopt a Prompt Learning approach integrated within the Prompt Spectrum. This approach utilizes specialized probabilistic embeddings that allow the model to handle uncertainty by generating prompts that represent a distribution of possible outcomes, rather than a single deterministic label. Our objective is to sample from the distribution defined by these prototype images.
@@ 3.3 Prompt-guided Diffusion Process
In this work, building on compositional prompts and the Prompt Spectrum, we extend the Stable Diffusion framework [29], a powerful latent diffusion model that operates in a compact latent space—dramatically reducing memory and computational demands compared to pixel-space diffusion— into the Progressive Spectrum Diffusion Model (PSDM). PSDM guides the image generation process through multi-granularity prompts $\mathbf{p}_{i}$, ensuring alignment with complex medical annotation requirements.
@@ 3.4 Training Strategy
Generating accurate images from diverse prompts while learning sequentially from different annotation types introduces a significant challenge known as catastrophic forgetting. This occurs when, as new annotations are introduced, the model ”forgets” how to generate images based on earlier annotations, diminishing its ability to retain previously learned information [31, 18]. While simultaneous training on all annotation modalities within a fixed dataset may appear intuitive, we observe that different modalities require varying numbers of training steps to converge effectively.
@@
$\underset{\theta}{\min}\mathcal{L}_{\text{Replay}}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,I),t,\mathbf{p}_{pre}\sim\mathcal{R}(\mathbf{p})}\left[\left\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{p}_{pre})\right\|^{2}\right]$
@@ 4.1.2 Downstream Tasks
Polyp Segmentation: For the segmentation task, following the PraNet standard, we trained our model on 1,450 image-mask pairs sourced from the Kvasir and CVC-ClinicDB datasets. Additionally, the PolypGen dataset (centers C1–C6), which serves as out-of-distribution data, was incorporated into the evaluation process to ensure a comprehensive assessment across diverse clinical environments. To improve the model’s generalization, we introduced two augmented datasets: the Segmentation Mask Only set, containing 1,450 image-mask pairs generated solely from the original training masks, and the Segmentation Mask-Text Description set, with 1,450 pairs generated using both textual descriptions and segmentation masks from the same training set.
@@
Polyp Detection: For detection, we used the PolypGen dataset(excluding 64 unlabeled C3 images). Training comprised 5,698 frames (1,385 C1–C5 images, 1,793 positive and 2,520 negative SequenceData frames from Seq1–15 and Seq1–13); testing comprised 2,275 frames (88 C6 images, 432 positive and 1,755 negative frames from Seq16–23 and Seq14–23). We also introduced two augmented datasets for detection: the Bounding Box Only set, with bounding boxes regressed from the Segmentation Mask Only set, and the Bounding Box-Text Description set, with bounding boxes from the Segmentation Mask-Text Description set. Bounding boxes were defined by the smallest rectangle enclosing each polyp in the segmentation masks.
@@ 4.1.3 Baseline Models and Implementations
For the segmentation task, we employ three state-of-the-art models: PraNet [42], Polyp-PVT [43], and Polyp-CASCADE [44]. PraNet was selected for its proven effectiveness in capturing both local and global features, Polyp-PVT for its superior performance in handling multi-scale features, and Polyp-CASCADE for its robust architecture in medical segmentation tasks. For the classification task, we adopt ResNet [45], a widely used model in medical image analysis known for its effectiveness and better suited than large-parameter Transformers such as ViT for small-sample experiments[46]. For the polyp detection task, we utilize YOLOv5 [47], chosen for its high performance and efficiency in real-time object detection, which is critical in clinical settings. The hyperparameters for each model are provided in Table 2, categorized by task. Each model was trained for 100 epochs, and the optimal values of the evaluation metrics were documented. Experiments for PSDM used Stable Diffusion 1.5 [29] with default settings: a DDIM sampler ($T=200$, guidance scale $w=7.5$) and the frozen CLIP encoder. The U-Net backbone consists of four down- and up-sampling stages, each containing two $3\times 3$ residual blocks, with a base channel width of 320 scaled by multipliers [1,2,4,4], and cross-attention layers at the middle two scales, totaling approximately 860 million parameters. All input images were resized to $256\times 256$px (mapped to a $64\times 64$ latent grid). Experiments ran on a single NVIDIA GeForce RTX 4090 (24 GB VRAM) under Ubuntu 20.04; fine-tuning required about 4 h of wall-clock time, with peak GPU memory usage around 24 GB. Our implementation uses Python 3.8.5 and CUDA 11.7.
4.1.4 Comparison with Other Generative Models
To comprehensively assess the performance of our proposed PSDM, we conduct comparisons with three representative generative models: the GAN-based Pix2Pix [48], the DDPM-based ArSDM [14], and the LDM-based Mask_C [26]. For fair and rigorous evaluation, all methods were trained under identical conditions, including the same dataset, random seeds, and classifier-guidance parameters, with an equal number of samples generated for quantitative analysis.
@@ 4.1.4 Comparison with Other Generative Models
Table 3: mDice and mIoU results for standard segmentation datasets. Best results are bold.
@@
Pixel to Pixel 88.83 81.57* 92.09* 86.67* 88.82* 83.04 73.67 65.74 62.84 56.37 75.47 68.42
ArSDM 86.13* 79.40* 90.66* 85.38* 87.92* 82.29* 72.71* 65.02* 68.91* 61.66* 76.08 69.02
Mask_C 88.81 81.58 90.91* 86.00* 89.86* 84.37* 72.24 65.01 57.38 57.38* 75.31 68.43
$\text{PSDM}^{\text{S}}$ 88.37 81.31 92.30* 87.60* 89.76* 84.12 74.00* 66.45* 63.55* 56.95* 75.91 69.09
$\text{PSDM}^{\text{M}}$ 88.49* 81.64* 92.08 86.99* 89.37 83.72 73.57* 66.21* 67.81* 60.98* 76.70 69.89
@@
Pixel to Pixel 87.45 79.49 91.71* 85.87* 91.01 85.40* 79.82* 71.28* 76.57* 67.52* 81.92 73.88
ArSDM 88.18* 80.54* 92.59* 87.51* 91.46* 86.47* 81.58 73.05 78.24* 70.08* 83.34 75.69
Mask_C 88.66* 81.26* 92.11* 86.57* 91.66 86.47 80.57 71.66* 77.12 68.12* 82.62 74.53
$\text{PSDM}^{\text{S}}$ 90.74* 84.33* 93.03* 88.14* 92.09* 87.40* 80.38 72.50 81.22* 73.13* 83.82 76.63
$\text{PSDM}^{\text{M}}$ 89.13 82.16 93.28* 88.39* 91.21* 86.12 81.23 73.06 78.73* 70.58* 83.40 75.96
@@
Pixel to Pixel 90.34* 83.54* 93.72* 89.12* 92.02 87.13* 81.50 73.34 79.53* 71.64* 83.94 76.64
ArSDM 88.99* 82.11* 93.13* 88.51* 90.64* 85.31* 82.29* 74.20* 79.94* 72.22* 84.10 76.81
Mask_C 89.49* 82.79* 93.78* 89.19* 92.11 87.11 82.00* 73.89* 77.89* 70.26 83.73 76.51
$\text{PSDM}^{\text{S}}$ 88.74* 82.01* 93.50 88.92 92.34 87.66 80.86 72.88* 79.33* 71.45* 83.50 76.31
$\text{PSDM}^{\text{M}}$ 89.26* 82.64* 94.02* 89.44* 92.01* 87.16* 82.28 73.93 80.31* 72.51* 84.45 77.10
Table 4: mDice and mIoU results for PolypGen. Best results are bold.
@@
Pixel to Pixel 78.70* 71.40* 72.79* 66.84* 87.03* 80.12* 36.85 31.65 53.36* 45.73* 73.43* 67.35* 70.10 63.52
ArSDM 81.13* 74.17 72.84 67.24 86.31 79.49* 36.17 31.01* 59.55* 50.97* 74.78 68.40* 71.12 64.55
Mask_C 79.47* 72.24* 73.20* 67.27* 86.87* 80.21* 37.69* 32.22 56.99 48.85 76.97 70.65 71.08 64.47
$\text{PSDM}^{\text{S}}$ 80.75* 73.92* 74.05* 68.31 87.51* 81.06 37.80* 32.52* 58.03* 50.08* 76.07* 70.01 71.76 65.38
$\text{PSDM}^{\text{M}}$ 81.89* 75.01 73.60 67.92* 87.35 80.78 39.52* 34.26* 60.80* 52.54* 76.70 70.25 72.48 66.01
@@
Pixel to Pixel 84.86* 77.25* 75.30* 69.46 88.35* 81.45 41.76 35.56 65.00 55.18* 77.42* 71.02* 74.55 67.47
ArSDM 84.89 77.55 76.66* 71.20 89.03 82.75 42.67 36.86 66.02 57.03 79.46* 73.61* 75.40 68.84
Mask_C 84.89* 77.50* 76.17* 70.57 88.93* 82.40* 43.20 37.12 65.40 56.33 77.76 71.12 75.36 68.45
$\text{PSDM}^{\text{S}}$ 85.11* 77.57* 75.84* 69.97* 88.76* 82.00* 44.28* 38.16* 67.11* 57.55* 76.56* 70.64* 75.42 68.47
$\text{PSDM}^{\text{M}}$ 84.77* 77.60* 76.23* 70.69* 88.93* 82.41* 43.86* 37.73* 68.03* 58.71* 76.69* 70.80* 75.56 68.84
@@
Pixel to Pixel 85.24* 78.63* 77.04 71.70 90.06* 84.02* 44.45 38.25 64.24* 55.30 78.25* 72.98* 75.82 69.43
ArSDM 85.97* 79.19* 76.32 71.00 90.01 83.88* 43.81 38.06* 66.06* 57.19* 80.52* 74.73* 76.04 69.67
Mask_C 86.08* 79.36* 76.48* 71.32* 89.82* 83.76 43.15* 37.43* 63.61* 54.95* 78.79 73.21 75.51 69.24
$\text{PSDM}^{\text{S}}$ 85.77* 79.12* 76.17* 70.85* 90.36* 84.37* 44.35 38.67* 67.13* 58.59* 78.23* 72.65* 76.18 69.94
$\text{PSDM}^{\text{M}}$ 86.17* 79.68* 76.37 71.11 89.71 83.64 45.25* 39.94* 68.04* 58.88* 77.67* 72.02 76.32 70.06
@@ 4.2.1 Polyp Segmentation
In our experiments, we evaluated the performance of the segmentation models using two key metrics: mean Dice coefficient (mDice) and mean Intersection over Union (mIoU). These metrics are crucial in medical image segmentation, as they measure the overlap between the predicted segmentations and the ground truth, with higher values indicating better segmentation performance. Statistical significance was evaluated using two-sided Wilcoxon signed-rank tests, with significant differences (p<<0.05) indicated by an asterisk (*) in the Table 3 and Table 4.
@@
Overall, these findings suggest that the proposed augmentation strategies consistently enhance models’ ability to generalize across both standard and challenging datasets. Compared with augmentations using polyp images from alternative generative models, our approach uniformly outperforms those augmentations across all evaluated segmentation architectures, yielding the highest overall mDice and mIoU on both standard benchmarks and the out‐of‐distribution PolypGen dataset. By integrating spatial and textual information through compositional prompts, the models show a superior capacity to handle variability in polyp characteristics. This improved generalization is crucial for real-world medical applications, where polyps exhibit significant heterogeneity in appearance, and accurate segmentation plays vital role in diagnostic precision.